OpenStudy (anonymous):

({x+1}/{x-1})=({-3}/{x+3})+({8}/{x^2+2x-3}) find X i factored, add but it seems i can't get the right answer. please help

OpenStudy (anonymous):

Solve for x over the real numbers: x^2+2 x-3 = 0 Factor the left hand side. The left hand side factors into a product with two terms: (x-1) (x+3) = 0 Solve each term in the product separately. Split into two equations: x-1 = 0 or x+3 = 0 Look at the first equation: Solve for x. Add 1 to both sides: x = 1 or x+3 = 0 Look at the second equation: Solve for x. Subtract 3 from both sides: Answer: | | x = 1 or x = -3

OpenStudy (anonymous):

Hopefully I helped!

OpenStudy (anonymous):

it marked me wrong, i guess the answer is neither 1 nor 0

OpenStudy (anonymous):

assuming what @zKrup did was right , neither 1 nor -3 are solution, cause they are zeroes of denominators, i.e there is no real solution